# Non-Iterative Disentangled Unitary Coupled-Cluster based on Lie-algebraic structure

### v2

@@ Abstract
Due to their non-iterative nature, fixed unitary coupled cluster (UCC) ansätze are attractive for performing quantum chemistry variational quantum eigensolver (VQE) computations as they avoid pre-circuit measurements on a quantum computer. However, achieving chemical accuracy for strongly correlated systems with UCC requires further inclusion of higher-order fermionic excitations beyond triples increasing circuit depth. We introduce $k$-NI-DUCC, a fixed and non-iterative disentangled unitary coupled cluster compact ansatz, based on specific ``$k^{\prime\prime}$ sets of “qubit” excitations, eliminating the needs for fermionic-type excitations. These elements scale linearly ($\mathcal{O}(n)$) by leveraging Lie algebraic structures, with $n$ being the number of qubits. The key excitations are screened through specific selection criteria, including the enforcement of all symmetries, to ensure the construction of a robust set of generators. NI-DUCC employs ``$k^{\prime\prime}$ products of the exponential of $\mathcal{O}(n)$- anti-Hermitian Pauli operators, where each single Pauli string has a length $p$. This results in a fewer two-qubit CNOT gates circuit scaling, $\mathcal{O}(knp)$, suitable for hardware implementations. Tested on LiH, H6 and BeH2, NI-DUCC-VQE achieves both chemical accuracy and rapid convergence even for molecules deviating significantly from equilibrium. It is hardware-efficient, reaching the exact full configuration interaction energy solution at specific layers, while reducing significantly the VQE optimization steps. If NI-DUCC-VQE effectively addresses the gradient measurement bottleneck of ADAPT-VQE-like iterative algorithms, the classical computational cost of constructing the $\mathcal{O}(n)$ set of excitations increases exponentially with the number of qubits. We provide a first implementation for constructing the generators’ set, able to handle up to 20 qubits and discuss the efficiency perspectives.
@@ Introduction
Therefore, a crucial element of VQE is the design of the parameterized ansatz. One initial approach in VQE relies on the use of the unitary coupled cluster singles and doubles (UCCSD) approach [22, 23, 24, 25]. It provides a well-defined fermionic excitation hierarchy, while maintaining size consistency and exhibiting variational properties. Furthermore, it achieves high accuracy and rapid convergence for near-equilibrium molecules, as demonstrated by successful experimental tests on small molecules [6, 26, 27, 28]. However, the non-commutativity among operators makes it challenging to directly translate the UCCSD ansatz into quantum gates, requiring the use of an additional Trotter approximation. In practice, this non-commutativity introduces complexities in operator ordering, which can significantly impact the resulting energy variance and potentially lead to non-uniqueness in the resultant wavefunction [29]. Moreover, as the molecular size increases, Trotterized UCCSD excitations may fail to achieve chemical accuracy while incorporating numerous unnecessary excitations. These limitations affect practical implementations on real hardware, as the increased circuit depth introduces additional noise [28]. Incorporating higher-order excitations remains possible to effectively capture the missing correlation effects of the Trotterized UCCSD ansatz. To address this, methods like Trotterized Symmetry unitary coupled cluster singles, doubles, triples (Sym-UCCSDT) and unitary selective coupled cluster singles, doubles, triples, quadruples (UsCCSDTQ) have been proposed. Sym-UCCSDT incorporates both spin and orbital symmetries [30], while UsCCSDTQ uses selected CI determinants [31]. Both methods aim to select only the necessary excitations, improving chemical accuracy and accelerating the optimization process in VQE for larger molecules. However, while they provide improvements upon the native Trotterized UCCSD ansatz, they also mechanically increase the circuit depth through the inclusion of higher-order excitations.
@@
To address these issues, constructing an ansatz tailored to system-specific criteria offers a promising solution for achieving shallower circuits. The popular ADAPT-VQE iterative algorithm is a notable example of this strategy [32]. While capable of extracting electronic correlation energy using a small subset of operators, the operator selection process corresponds itself to a high computational overhead associated to gradients measurements. Despite subsequent efforts to minimize these costs, ADAPT-VQE versions [33, 34, 35, 36, 37, 38, 39, 40] continue to face gradient measurement challenges, especially as molecular size increases, while maintaining the necessary chemical accuracy. Similar to ADAPT-VQE methods, the Qubit coupled cluster (QCC) approach [41, 42] and its variants, including iterative Qubit coupled cluster (iQCC) [43], offer promising strategies for designing a compact wavefunction. In iQCC, generators are chosen from a specific set of operators, including all possible Pauli strings, to ensure energy reduction. These generators are sampled using first-order derivative calculations, resulting in a shallow circuit depth. Unlike ADAPT-VQE, the iQCC method avoids the need for full re-optimization of all parameters at each iteration, enhancing its practicality for NISQ devices. However, this advantage comes with costs: each iteration requires a canonical transformation of the Hamiltonian, which leads to an expanded Hamiltonian size. In practice, this requires high computational resources on classical computers, coupled to additional measurements on quantum devices, and gradient calculations at each step. Furthermore, the iQCC method does not fully guarantee the ordering of operators[43], which can increase the risk of encountering local minima and contribute to the ‘barren plateaus’ problem [44]. Nevertheless, recent advances have been made in exploring efficient techniques for QCC methods, such as in [45, 46]. In addition to the qubit coupled cluster (QCC) and ADAPT-VQE approaches, recent advancements such as the tiled unitary product state (tUPS) [47] and subspace projective quantum eigensolver (SPQE) [48] methods have significantly improved computational accuracy and the efficiency of quantum computing resources.
@@
In this paper, we introduce the non-iterative disentangled unitary coupled cluster (NI-DUCC) method, a novel and efficient approach designed for NISQ devices. NI-DUCC simplifies the unitary coupled cluster framework by operating directly in qubit space excitations, that are optimally generated through Lie algebraic structures. These operators (each of length $p$) are anti-Hermitian and their set scales linearly with the number of qubits, $n$, which significantly reduces the reliance on error-prone two-qubit CNOT gates (see “Quantum computer datasheet”, Google, 2021 [49]). The NI-DUCC-VQE algorithm avoids gradient measurements and integrates symmetries to accelerate convergence and optimization. Additionally, it leverages Lie algebraic relations to resolve the ordering issues commonly encountered in standard UCC methods that rely on Trotterization. Our numerical simulations across various molecules show that NI-DUCC effectively reduces VQE energy without barren plateaus problems. It achieves Full Configuration Interaction (FCI) solutions as the number of layers ($k$) increases and it scales CNOT gate counts as $\mathcal{O}(knp)$ per circuit, which makes it a promising approach for current quantum hardware.
@@
The rest of the paper is organized as follows. The first section provides an introduction to the NI-DUCC ansatz, followed by a detailed explanation of the NI-DUCC-VQE algorithm. In Section 2, we describe the computational methods and present various numerical tests on several molecules. To showcase the NI-DUCC-VQE capabilities, especially in achieving chemical accuracy, we compare its results to those obtained from UCCSD, UCCSDT, ADAPT-VQE, UsCCSDTQ, and to the reference FCI classical approach. In the final section, we examine the NI-DUCC resource requirements and compare its performance to that of current state-of-the-art methods.
@@ I.1 Disentangled unitary coupled cluster (DUCC): Lie Algebraic structure and symmetric excitations
with $\hat{P}_{i},\hat{P}_{j},\hat{P}_{r}\in\mathcal{S}$. In practice, utilizing a set $\mathcal{S}$ that grows exponentially with $n$ is impractical, as it would compromise the efficiency of using quantum computers to represent $\hat{U}(\vec{\theta})$. Instead, we aim to reduce the set of excitations into a size that corresponds to a linear function of $n$. To do so, we leverage the advantage of retaining the set of starters $\mathcal{S}_{j}$, as they effectively help in lowering energy and enable a faster optimization. Then, it requires to find a set $\mathcal{S}_{j+1}$ with an optimal number of $\hat{P}_{l}$ such that the union $\mathcal{S}^{(c)}=\big{\{}\mathcal{S}_{j}\cup\mathcal{S}_{j+1}\big{\}}$ generates a Lie sub-algebra of length $\ll 4^{n-1}$. To determine the optimal size of $\mathcal{S}_{j+1}$, we reviewed the findings from ref. [37]. It was demonstrated that a minimum complete pool (MCP) of $22n-$ 2 Pauli operators strings is sufficient for achieving chemical accuracy with qubit ADAPT-VQE for molecules such as H4, LiH, and BeH2. Following the guidelines for constructing an MCP proposed in reference [37], we generate excitations in the NI-DUCC approach with the set $\mathcal{S}^{(c)}=\big{\{}\mathcal{S}_{j}\cup\mathcal{S}_{j+1}\big{\}}$, ensuring that its size is $22n-$ 2:
@@
As a result, the above equation represents a “compact” and “ordered” DUCC wavefunction, relying heavily on a sufficient number of starters, $\mathcal{S}_{j}$, that adhere to all symmetries. However, the “strength” of these symmetric starters significantly influences convergence, as demonstrated in our simulations (see Section (III.1)). This strength is measured using fermionic pre-screening excitations, as detailed in the algorithm below. Given that the NI-DUCC wavefunction operates directly within the space of multi-qubit operators, $\hat{P_{l}}$, with minimal complexity of two-qubit CNOT gates, it can be tailored to fit specific hardware constraints. In this line, the NI-DUCC approach can be connected to the hardware efficient ansatz [54], which employs multiple layers of symmetry-preserving gates with various variational parameters. Thus we can similarly implement the NI-DUCC wavefunction by incorporating $k$ layers, each consisting of of $22n-$ 2 operators.
@@
This wavefunction offers flexibility as it allows users to adjust the depth of layers $k$ in the circuit. As shown in the results section, increasing the number of layers improves accuracy. If a certain level of layers is added, it even allows to reach the FCI energy solution, with fewer optimization steps. Moreover, the NI-DUCC approach avoids the need for iterative energy gradients, which is a primary bottleneck of ADAPT-VQE (see Table S1 in the supplementary materials [55] for the scaling of residual gradients in various ADAPT-VQE versions).
@@ I.2 The NI-DUCC-VQE protocol
In this section, we discuss the 4 preparatory components of NI-DUCC-VQE algorithm. We refer the reader to section 2 of the supplementary materials [55], which systematically describes the basic outline of the algorithm.
Figure 1: Schematic of NI-DUCC-VQE algorithm. (1) The necessary conditions for constructing a robust symmetry-preserving minimal complete pool (MCP) of size $22n-$ 2. The initial qubit excitations (starters) are strong and symmetric, effectively screened from fermionic excitations. (2) Once the MCP is generated, the NI-DUCC ansatz is constructed from layers $k$. To the left, an example of mapping the exponential of a Pauli operator into a circuit is shown, by using staircase method [7, 10, 20]. (3) Numerical results tested on BeH2 (14 qubits), using NI-DUCC-VQE, showing the comparison between robust and non-robust MCP excitations, in terms of chemical accuracy. (4) A list of applications that can benefit from NI-DUCC-VQE and its ansatz.
First, we compute the one- and two-electron integrals denoted as $h_{ij}$ (Eq. 2 of the supplementary material [55]) and $v_{ijkl}$ (Eq. 3 of the supplementary material [55]), which is a task that can be efficiently performed on a classical computer within polynomial time. Subsequently, we calculate the electronic structure Hamiltonian (Eq. 1 of the supplementary material [55]), a standard step in every VQE algorithm. Second, we generate all double excitations required to construct the selective UsCCD ansatz, which is a simplified version of UsCCSDTQ [31]. We denote these excitations by the $(i,j,k,l)$ indices and add them to a list, denoted $D$. We then initialize an array of zeros for the parameters associated with these double excitation operators, which serve as the initial guess for the UsCCD-VQE process. For the third component, we select the Hartree-Fock state $|\psi_{HF}\rangle$, as the initial reference state.
Then our objective is to construct the unitary operator described in Eq. (3), thereby completing the NI-DUCC wavefunction (Eq. 4). To achieve this, we will outline the following three steps:

### v3

@@ Abstract
Due to their non-iterative nature, fixed unitary coupled cluster (UCC) ansätze are attractive for performing quantum chemistry variational quantum eigensolver (VQE) computations as they avoid pre-circuit measurements on a quantum computer. However, achieving chemical accuracy for strongly correlated systems with UCC requires further inclusion of higher-order fermionic excitations beyond triples increasing circuit depth. We introduce $k$-NI-DUCC, a fixed and non-iterative disentangled unitary coupled cluster compact ansatz, based on specific “$k$” sets of “qubit” excitations, eliminating the needs for fermionic-type excitations. These elements scale linearly ($\mathcal{O}(n)$) by leveraging Lie algebraic structures, with $n$ being the number of qubits. The key excitations are screened through specific selection criteria, including the enforcement of all symmetries, to ensure the construction of a robust set of generators. NI-DUCC employs “$k$” products of the exponential of $\mathcal{O}(n)$- anti-Hermitian Pauli operators, where each single Pauli string has a length $p$. This results in a fewer two-qubit CNOT gates circuit scaling, $\mathcal{O}(knp)$, suitable for hardware implementations. Tested on LiH, H6 and BeH2, NI-DUCC-VQE achieves both chemical accuracy and rapid convergence even for molecules deviating significantly from equilibrium. It is hardware-efficient, reaching the exact full configuration interaction energy solution at specific layers, while reducing significantly the VQE optimization steps. While NI-DUCC-VQE effectively addresses the gradient measurement bottleneck of ADAPT-VQE-like iterative algorithms, the classical computational cost of constructing the $\mathcal{O}(n)$ set of excitations increases exponentially with the number of qubits. We provide a first implementation for constructing the generators’ set, able to handle up to 20 qubits and discuss the efficiency perspectives.
@@ Introduction
Therefore, a crucial element of VQE is the design of the parameterized ansatz. One initial approach in VQE relies on the use of the unitary coupled cluster singles and doubles (UCCSD) approach [22, 23, 24, 25]. It provides a well-defined fermionic excitation hierarchy, while maintaining size consistency and exhibiting variational properties. Furthermore, it achieves high accuracy and rapid convergence for near-equilibrium molecules, as demonstrated by successful experimental tests on small molecules [6, 26, 27, 28]. However, the non-commutativity among operators makes it challenging to directly translate the UCCSD ansatz into quantum gates, requiring the use of an additional Trotter approximation. In practice, this non-commutativity introduces complexities in operator ordering, which can significantly impact the resulting energy variance and potentially lead to non-uniqueness in the resultant wavefunction [29]. Moreover, as the molecular size increases, Trotterized UCCSD excitations may fail to achieve chemical accuracy while incorporating numerous unnecessary excitations. These limitations affect practical implementations on real hardware, as the increased circuit depth introduces additional noise [28]. Incorporating higher-order excitations remains possible to effectively capture the missing correlation effects of the Trotterized UCCSD ansatz. To address this, methods such as Trotterized Symmetry unitary coupled cluster singles, doubles, triples (Sym-UCCSDT) and unitary selective coupled cluster singles, doubles, triples, quadruples (UsCCSDTQ) have been proposed. Sym-UCCSDT incorporates both spin and orbital symmetries [30], while UsCCSDTQ uses selected CI determinants [31]. Both methods aim to select only the necessary excitations, improving chemical accuracy and accelerating the optimization process in VQE for larger molecules. However, while they provide improvements upon the native Trotterized UCCSD ansatz, they also mechanically increase the circuit depth through the inclusion of higher-order excitations.
@@
To address these issues, constructing an ansatz tailored to system-specific criteria offers a promising solution for achieving shallower circuits. The popular ADAPT-VQE iterative algorithm is a notable example of this strategy [32]. It effectively extracts electronic correlation energy using a small set of operators. However, its operator selection results in high computational overhead due to gradient measurements. These measurements are essential for evaluating the impact of adding candidate operators to the ansatz. At each iteration, the algorithm calculates the energy gradient, indicating how much the energy changes with each addition. This process is expensive and leads to increasing overhead as the ansatz grows. Despite subsequent efforts to minimize these costs, ADAPT-VQE variants [33, 34, 35, 36, 37, 38, 39, 40] continue to face gradient measurement challenges, especially as molecular size increases, while maintaining the necessary chemical accuracy. Similar to ADAPT-VQE methods, the Qubit coupled cluster (QCC) approach [41, 42] and its variants, including iterative Qubit coupled cluster (iQCC) [43], offer promising strategies for designing a compact wavefunction. In iQCC, generators are chosen from a specific set of operators, including all possible Pauli strings, to ensure energy reduction. These generators are sampled using first-order derivative calculations, resulting in a shallow circuit depth. Unlike ADAPT-VQE, the iQCC method avoids the need for full re-optimization of all parameters at each iteration, enhancing its practicality for NISQ devices. However, this advantage comes with costs: each iteration requires a canonical transformation of the Hamiltonian, which leads to an expanded Hamiltonian size. In practice, this requires high computational resources on classical computers, coupled to additional measurements on quantum devices, and gradient calculations at each step. Furthermore, the iQCC method does not fully guarantee the ordering of operators[43], which can increase the risk of encountering local minima and contribute to the ‘barren plateaus’ problem [44]. Nevertheless, recent advances have been made in exploring efficient techniques for QCC methods, such as in [45, 46]. In addition to the qubit coupled cluster (QCC) and ADAPT-VQE approaches, recent advancements such as the tiled unitary product state (tUPS) [47], the COMmutativity Pre-screened Automated Selection of Scatterers (COMPASS) [48], the Cognitive Optimization via Many-body Perturbation theory towards Ansatz Construction and Tailoring (COMPACT) [49], and subspace projective quantum eigensolver (SPQE) [50] methods have significantly improved computational accuracy and the efficiency of quantum computing resources.
@@
In this paper, we introduce the non-iterative disentangled unitary coupled cluster (NI-DUCC) method, a novel and efficient approach designed for NISQ devices. NI-DUCC simplifies the unitary coupled cluster framework by operating directly in qubit space excitations, that are optimally generated through Lie algebraic structures. These operators (each of length $p$) are anti-Hermitian and their set scales linearly with the number of qubits, $n$, which significantly reduces the reliance on error-prone two-qubit CNOT gates (see “Quantum computer datasheet”, Google, 2021 [51]). The NI-DUCC-VQE algorithm avoids gradient measurements and integrates symmetries to accelerate convergence and optimization. Additionally, it leverages Lie algebraic relations to overcome ordering issues typically seen in Trotterized UCC methods. This may explain why our simulations across various molecules show that NI-DUCC effectively reduces VQE energy without barren plateau issues [52]. While also ADAPT-VQE is considered relatively immune to barren plateaus and local minima as shown in some simulations [53], it still faces challenges with gradient troughs that can hinder smooth convergence. In contrast, NI-DUCC achieves consistent VQE energy reduction, offering a smoother and more reliable convergence path. Furthermore, it achieves Full Configuration Interaction (FCI) solutions as the number of layers ($k$) increases, requiring significantly fewer optimization steps, and it scales CNOT gate counts as $\mathcal{O}(knp)$ per circuit, which makes it a promising approach for current quantum hardware.
@@
The rest of the paper is organized as follows. The first section provides an introduction to the NI-DUCC ansatz, followed by a detailed explanation of the NI-DUCC-VQE algorithm. In Section 2, we describe the computational methods and present various numerical tests on several molecules. To showcase the NI-DUCC-VQE capabilities, especially in achieving chemical accuracy, we compare its results to those obtained from UCCSD, UCCSDT, sym-UCCSDT, UsCCSDTQ, COMPASS, COMPACT, ADAPT-VQE, and to the reference FCI classical approach. In the final section, we examine the NI-DUCC resource requirements and compare its performance to that of current state-of-the-art methods.
@@ I.1 Disentangled unitary coupled cluster (DUCC): Lie Algebraic structure and symmetric excitations
with $\hat{P}_{i},\hat{P}_{j},\hat{P}_{r}\in\mathcal{S}$. In practice, utilizing a set $\mathcal{S}$ that grows exponentially with $n$ is impractical, as it would compromise the efficiency of using quantum computers to represent $\hat{U}(\vec{\theta})$. Instead, we aim to reduce the set of excitations into a size that corresponds to a linear function of $n$. To do so, we leverage the advantage of retaining the set of starters $\mathcal{S}_{j}$, as they effectively help in lowering the energy and enable a faster optimization. Then, it requires to find a set $\mathcal{S}_{j+1}$ with an optimal number of $\hat{P}_{l}$ such that the union $\mathcal{S}^{(c)}=\big{\{}\mathcal{S}_{j}\cup\mathcal{S}_{j+1}\big{\}}$ generates a Lie sub-algebra of length $\ll 4^{n-1}$. To determine the optimal size of $\mathcal{S}_{j+1}$, we reviewed the findings from reference [37]. It was demonstrated that a minimum complete pool (MCP) of $22n-$ 2 Pauli operators strings is sufficient for achieving chemical accuracy with qubit ADAPT-VQE for molecules such as H4, LiH, and BeH2. Following the guidelines for constructing an MCP proposed in reference [37], we generate excitations in the NI-DUCC approach with the set $\mathcal{S}^{(c)}=\big{\{}\mathcal{S}_{j}\cup\mathcal{S}_{j+1}\big{\}}$, ensuring that its size is $22n-$ 2:
@@
As a result, the above equation represents a “compact” and “ordered” DUCC wavefunction, relying heavily on a sufficient number of starters, $\mathcal{S}_{j}$, that adhere to all symmetries. However, the “strength” of these symmetric starters significantly influences convergence, as demonstrated in our simulations (see Section (III.1)). This strength is measured using fermionic pre-screening excitations, as detailed in the algorithm below. Given that the NI-DUCC wavefunction operates directly within the space of multi-qubit operators, $\hat{P_{l}}$, with minimal complexity of two-qubit CNOT gates, it can be tailored to fit specific hardware constraints. In this line, the NI-DUCC approach can be connected to the hardware efficient ansatz [58], which employs multiple layers of symmetry-preserving gates with various variational parameters. Thus we can similarly implement the NI-DUCC wavefunction by incorporating $k$ layers, each consisting of $22n-$ 2 operators (see step 2 in Figure 1).
@@
This wavefunction offers flexibility as it allows users to adjust the depth of layers $k$ in the circuit. As shown in the results section, increasing the number of layers improves accuracy. If a certain level of layers is added, it even allows to reach the FCI energy solution, with fewer optimization steps. For the systems studied (detailed in Section IV), k-NI-DUCC consistently reached FCI solution with $k=8$ layers, demonstrating efficient scaling in capturing correlations. We found that the NI-DUCC ansatz outperformed existing methods in CNOT gate efficiency but it requires 2-3 times more parameters than ADAPT-VQE, depending on the system. In terms of general applicability, while $k$ can sometimes be roughly estimated based on system size and bond length deviations, it is typically more reliable to iteratively test performance across values of $k$ to identify the optimal layer count for a given accuracy. This iterative approach helps ensure efficient resource use in terms of parameters and CNOT gates, particularly for larger molecular systems where expressivity demands can significantly vary. In addition to its flexible structure, NI-DUCC bypasses the need for iterative energy gradient calculations, a major bottleneck in ADAPT-VQE (we refer the reader to Table S1 in the supplementary materials [59] for the scaling of CNOT gates, Parameter counts and residual gradients across different versions of ADAPT-VQE).
@@ I.2 The NI-DUCC-VQE protocol
In this section, we discuss the 4 preparatory components of NI-DUCC-VQE algorithm. We refer the reader to section 2 of the Supplementary Materials [59], which systematically describes the basic outline of the algorithm. First, we compute the one- and two-electron integrals denoted as $h_{ij}$ (Eq. 2 of the supplementary material [59]) and $v_{ijkl}$ (Eq. 3 of the Supplementary Materials [59]), which is a task that can be efficiently performed on a classical computer within polynomial time. Subsequently, we calculate the electronic structure Hamiltonian (Eq. 1 of the Supplementary Materials [59]), a standard step in every VQE algorithm. Second, we generate all double excitations required to construct the selective UsCCD ansatz, which is a simplified version of UsCCSDTQ [31]. We denote these excitations by the $(i,j,k,l)$ indices and add them to a list, denoted $D$. We then initialize an array of zeros for the parameters associated with these double excitation operators, which serve as the initial guess for the UsCCD-VQE process. For the third component, we select the Hartree-Fock state $|\psi_{HF}\rangle$, as the initial reference state. Then our objective is to construct the unitary operator described in Eq. (3), thereby completing the NI-DUCC wavefunction (Eq. 4). To accomplish this, we will outline the following three steps, each visualized in the schematic of the NI-DUCC-VQE algorithm in Figure 1 below:
@@
Figure 1: Schematic of NI-DUCC-VQE algorithm. (1) The necessary conditions for constructing a robust symmetry-preserving minimal complete pool (MCP) of size $22n-$ 2. (1.c) The initial qubit excitations (starters) are strong, symmetric operators, selected via a fermionic pre-screening criterion based on a threshold $\epsilon$, ensuring only dominant excitations are included while weaker excitations are filtered out. (1.d) Verifying the completeness of the set of the chosen starters (More details can be found in section 4 of the Supplementary Materials
[59]). (2) Once the MCP is generated, the NI-DUCC ansatz is constructed from $k$ layers. To the left, an example of mapping the exponential of a Pauli operator into a circuit is shown, by using staircase method [7]. (3) Numerical results tested on BeH2 (14 qubits), using NI-DUCC-VQE, showing the comparison between robust and non-robust MCP excitations, in terms of chemical accuracy. (4) A list of applications that can benefit from NI-DUCC-VQE and its ansatz. For further details on extending the NI-DUCC method to excited states or open-shell molecules, including techniques for carefully enforcing symmetries, please refer to Section 4.A in the Supplementary Materials[62].